X Company is considering buying a part in 2020 that it made in 2019. For 2020, the budgeted production cost function is $10.43X + $9,000, where X is the number of units produced. A company has offered to supply this part to X Company for $13.66 per unit. If X Company accepts the offer, it will still incur fixed costs of $4,770, and it will not be able to do anything with the resources that were used to produce the part. In 2019, 3,000 units of the part were produced; in 2020, 3,400 units will be needed.

1. If X Company makes the part in 2020 instead of buying it, it will save?

2. At what production level would X Company be indifferent between making and buying the part in 2020?
Answer #1

1.

Total cost of making 3,400 units in 2020 = 10.43X + 9,000

= 10.43 x 3,400 + 9,000

= 35,462+9,000

= $44,462

Total cost of buying 3,400 units in 2020 = 13.66 X + 4,770

= 13.66 x 3,500 + 4,770

= 46,444+4,770

= $51,214

If X company makes the part in 2020 instead of buying it, it will save = Total cost of buying 3,400 units in 2020 - Total cost of making 3,400 units in 2020

= 51,214-44,462

= $6,752

2.

Let the production level of K units, total cost of making and buying is same.

Total cost of making = Total cost of buying

10.43K + 9,000 = 13.66K+4,770

13.66K-10.43K = 9,000-4,770

3.23K = 4,230

K = 1,310

Hence, X company will be indifferent between making and buying the part it production level of 1,310 units.

Kindly comment if you need further assistance. Thanks‼!